OpenStudy (jaedyn):

Write an expression that is equivalent to three over four (5z+ 16). 15 over fourz + 16 15 over fourz + 12 three over fourz + 16 three over fourz + 12

OpenStudy (jaedyn):

Help

OpenStudy (3mar):

May I ?

OpenStudy (jaedyn):

Yes

OpenStudy (3mar):

thanks

OpenStudy (3mar):

What did you pick?

OpenStudy (3mar):

\[\frac{ 3 }{ 4 }(5z+16)=\frac{ 3 }{ 4 }*5z+\frac{ 3 }{ 4 }*16=\frac{ 15z }{ 4 }+12\]

OpenStudy (evansanchez):

3/4(5z+16)

OpenStudy (3mar):

Thank you for the medal!